# DAL: A Practical Prior-Free Black-Box Framework for Non-Stationary Bandit Environments

### v2

@@ B.1 Regret Bounds of DAL in Piecewise Stationary Environments
As discussed in Section 4.2 of the paper, using Corollary 4.4, we can select different algorithms as input for DAL to attain or improve the state-of-the-art regret bounds in PS environments. Combining DAL with different bandit algorithms leads to the results of Table 1. It is evident that DAL matches the state-of-the-art regret bounds in PS-LBs and PS-GLBs, and DAL improves the best known bounds in the PS-SCB and PS-KB settings. Note that for PS-SCBs, the strongest result correspond to the prior-based WeightUCB [57]. As demonstrated in the final columns of the table, the order-wise dependence on problem parameters from the stationary setting seamlessly transfers to the PS setting without degradation.
@@
PS Setting Non-Stationary Algorithm NS Algorithm Regret Bound in $\tilde{\mathcal{O}}(\cdot)$ DAL Input Regret Bound in $\tilde{\mathcal{O}}(\cdot)$
PS-LB MASTER [58] + LinUCB $d\sqrt{TN_{T}}$
ADA-OPKB [26] $d\sqrt{N_{T}T}$
@@
PS-KB MASTER [58] + GPUCB $\gamma_{T}\sqrt{N_{T}T}$
ADA-OPKB [26] $\sqrt{d\gamma_{T}N_{T}T}$
@@
In this subsection, we establish the construction of $\mathcal{A}_{e}$ in the NS-KB setting. According to Lemma 5 from [15], we have that every $f\in H_{k}$ with $\|f\|_{H_{k}}\leq B$ is Lipschitz continuous, satisfying the following,
@@
Recall that $\mathcal{P}_{T}$ corresponds to the set of centers of the balls of an arbitrary $\delta_{T}$-cover of $\mathcal{A}\subseteq[0,R]^{d}$. Let $[a]_{\mathrm{e}}$ denote the action in $\mathcal{P}_{T}$ that is the closest to $a\in\mathcal{A}$, i.e., $[a]_{\mathrm{e}}=\mathrm{argmin}_{x\in\mathcal{P}_{T}}\|a-x\|_{2}$. Then, we can leverage the Lipschitz property of functions in the RKHS to obtain the following upper bound: For any $a\in\mathcal{A}$ and $f\in H_{k}$ with $\|f\|_{H_{k}}\leq B$,
@@
$\overset{(a)}{\geq}|f(\tilde{a})-f^{\prime}(\tilde{a})|-2BL_{u}\delta_{T}$
>$\overset{(b)}{>}0$
where step $(a)$ is due to (1), and step $(b)$ is due to the assumption in Proposition 4.2. This indicates that the value of the reward function at $[\tilde{a}]_{\mathrm{e}}$ must change by a non-zero amount. Thus, one can use observations from action $[\tilde{a}]_{\mathrm{e}}$ in order to deduce whether the reward function has changed its value in action $\tilde{a}$. In addition, by the upper bound on the covering number, the cardinality of $\mathcal{P}_{T}$ is upper bounded by $\lceil\sqrt{d}R/2\delta_{T}\rceil^{d}$.
@@
The proof of Theorem 4.3 follows exactly those of Theorem 1 and Corollary 1 in [27], with $A$ replaced by $N_{\mathrm{e}}$, due to the different number of actions in the covering set. For completeness, we provide a proof sketch of Theorem 4.3: First, we partition the regret into two cases. If no false alarm occurs and all changes are detected within a short delay, we can separate the regret into three components: the regret due to forced exploration, the regret during the short detection (restart) delay after changes, and the regret incurred by the stationary bandit algorithm after the change is detected. If not, we use a crude linear bound to bound the regret and show that the probability of false alarm and that of late detection are low, which ensures that the regret due to detection failure is small.
@@
In PS-PBs, $N_{\mathrm{e}}=d$, $p\geq 1/2$, and $q=0$. Thus, $R_{T}=\tilde{\mathcal{O}}(\sqrt{dN_{T}T}+d^{p}\sqrt{N_{T}T})=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{N_{T}T})$.
In PS-KBs, $p,q\geq 1/2$. We can upper bound $N_{\mathrm{e}}$ using the fact that $|\mathcal{P}_{T}|\leq\lceil\sqrt{d}R/2\delta_{T}\rceil^{d}$. Thus, $N_{e}\leq\lceil C\gamma_{T}^{2q/d}\rceil^{d}$ with $\delta_{T}=\frac{Rd^{1/2-2p/d}}{2(C\gamma_{T}^{2q})^{1/d}}$ and $R_{T}=\tilde{\mathcal{O}}(\sqrt{d^{2p}\gamma_{T}^{2q}N_{T}T}+d^{p}\gamma_{T}^{q}\sqrt{N_{T}T})=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{N_{T}T})$.
We emphasize that when the number of action is smaller than the covering number, i.e., $|\mathcal{A}|<\lceil C\gamma_{T}^{2q/d}\rceil^{d}\leq\gamma_{T}$, then we can set $\mathcal{A}_{e}$ to be the entire action set $\mathcal{A}$. In this case, $N_{\mathrm{e}}<\gamma_{T}$, guaranteeing order-optimal regret.

### v4

@@ B.1 Regret Bounds of DAL in Piecewise Stationary Environments
As discussed in Section 4.2 of the paper, using Corollary 4.5, we can select different algorithms as input for DAL to attain or improve the state-of-the-art regret bounds in PS environments. Combining DAL with different bandit algorithms leads to the results in Table 1. It is evident that DAL matches the state-of-the-art regret bounds in PS-LBs, PS-GLBs and PS-CBs, and DAL improves the best known bounds in the PS-SCB and PS-KB settings. Note that for PS-SCBs, the strongest result corresponds to the prior-based WeightUCB [61]. As demonstrated in the final columns of the table, the order-wise dependence on problem parameters from the stationary setting seamlessly transfers to the PS setting without degradation.
@@
[…]
PS-LB MASTER [62] + LinUCB $d\sqrt{TN_{T}}$ -
ADA-OPKB [30] $d\sqrt{N_{T}T}$ -
@@
PS-KB MASTER [62] + GPUCB $\gamma_{T}\sqrt{N_{T}T}$ -
ADA-OPKB [30] $\sqrt{d\gamma_{T}N_{T}T}$ -
@@
PS-CB MASTER [62] + ILTCB $\sqrt{|\mathcal{A}|N_{T}T\log|\Pi|}$ -
ADA-ILTCB+ [15] $\sqrt{|\mathcal{A}|N_{T}T\log|\Pi|}$ -
DAL (ours) + ILTCB [2] $\sqrt{|\mathcal{A}|N_{T}T\log|\Pi|}$ $\sqrt{|\mathcal{A}|T\log|\Pi|}$
DAL (ours) + SquareCB [26] $\sqrt{|\mathcal{A}|N_{T}T\log|\Pi|}$ $\sqrt{|\mathcal{A}|T\log|\Pi|}$
@@
In this subsection, we establish the construction of $\mathcal{A}_{\mathrm{e}}$ in the NS-KB setting. According to Lemma 5 from [18], we have that every $f\in H_{k}$ with $\|f\|_{H_{k}}\leq B$ is Lipschitz continuous, satisfying the following,
@@
Recall that $\mathcal{V}_{T}$ corresponds to the set of centers of the balls of an arbitrary $\delta_{T}$-cover of $\mathcal{A}\subseteq[0,R]^{d}$. Let $a]_{\mathrm{e}}$ denote the action in $\mathcal{V}_{T}$ that is the closest to $a\in\mathcal{A}$, i.e., $a]_{\mathrm{e}}=\mathrm{argmin}_{x\in\mathcal{P}_{T}}\|a-x\|_{2}$. Then, we can leverage the Lipschitz property of functions in the RKHS to obtain the following upper bound: For any $a\in\mathcal{A}$ and $f\in H_{k}$ with $\|f\|_{H_{k}}\leq B$,
@@
$\overset{(a)}{\geq}|f(\tilde{a})-f^{\prime}(\tilde{a})|-2BL_{u}\delta_{T}\overset{(b)}{>}0$
where step $(a)$ is due to (1), and step $(b)$ is due to the assumption in Proposition 4.2. This indicates that the value of the reward function at $\tilde{a}]_{\mathrm{e}}$ must change by a non‑zero amount. Thus, one can use observations from action $\tilde{a}]_{\mathrm{e}}$ in order to deduce whether the reward function has changed its value in action $\tilde{a}$. In addition, by the upper bound on the covering number, the cardinality of $\mathcal{V}_{T}$ is upper bounded by $\lceil\sqrt{d}R/2\delta_{T}\rceil^{d}$.
@@
For PS-PBs and PS-KBs, the proof of Theorem 4.3 follows exactly the same as those of Theorem 1 and Corollary 1 in [31], with the number of arms replaced by $N_{\mathrm{e}}$, due to the different number of actions in the covering set. For completeness, we provide a proof sketch of Theorem 4.3: First, we partition the regret into two cases. If no false alarm occurs and all changes are detected within a short delay, we can separate the regret into three components: the regret due to forced exploration, the regret during the short detection (restart) delay after changes, and the regret incurred by the stationary bandit algorithm after the change is detected. If not, we use a crude linear bound to bound the regret and show that the probability of false alarm and that of late detection are low, which ensures that the regret due to detection failure is small.
For PS-CBs, the proof of Theorem 4.3 follows similar to those of NS-KBs and NS-PBs, except that the definition of successful detection events should be modified as follows: For any $k\in[N_{T}]$,
$\mathcal{M}_{k}\coloneqq$
$\Bigg\{\!\forall l\in[k-1],c\in\mathcal{C},i\in[N_{\mathrm{e}}]:\!\sum_{t=\tau_{l-1}}^{\nu_{l}-1}\!\mathds{1}\bigg\{\!C_{t}=c,(t-\tau_{l-1}-1)\!\!\!\!\!\mod\!\!\left\lceil\frac{N_{\mathrm{e}}}{\alpha_{l}}\right\rceil\!=i\!-\!1\!\bigg\}\!\geq m_{\mathcal{D}}\!\Bigg\},$ (2)
$\mathcal{L}_{k}\coloneqq$
$\Bigg\{\!\forall l\in[k-1],c\in\mathcal{C},i\in[N_{\mathrm{e}}]:\!\sum_{t=\nu_{l}}^{\nu_{l}+\ell_{l}-1}\!\mathds{1}\bigg\{\!C_{t}=c,(t-\tau_{l-1}-1)\!\!\!\!\!\mod\!\!\left\lceil\frac{N_{\mathrm{e}}}{\alpha_{l}}\right\rceil\!=i\!-\!1\!\bigg\}\!\geq\ell_{\mathcal{D}}\!\Bigg\},$ (3)
$\mathcal{E}_{k}\coloneqq\left\{\forall\,l\in[k-1]:\tau_{l}\in\left\{\nu_{l},\dots,\nu_{l}+\ell_{l}-1\right\}\right\}\cap\mathcal{M}_{k}\cap\mathcal{L}_{k},\>\mathrm{and}\>$ (4)
$\mathcal{G}_{k}\coloneqq\mathcal{E}_{k}\cap\left\{\tau_{k}\geq\nu_{k}\right\},$ (5)
where $\tau_{l}$ denotes the $l^{\mathrm{th}}$ detection point, and recall that $\ell_{l}=\lceil N_{\mathrm{e}}/\alpha_{l}\rceil[2\log(T)+\ell_{\mathcal{D}}/s]$ with $s=\min_{c\in\mathcal{C},t\in[T]:\mathcal{P}_{t}(c)>0}\mathcal{P}_{t}(c)$ for $l\in[N_{T}]$. For convenience in notations, we denote $\tau_{0}=0$ and $0]=\emptyset$. In the new definition of the detection success event, the number of reward samples for each context-action pair obtained from forced exploration is lower bounded, ensuring enough pre- and post-change samples for change detection. With these new successful detection events, the proof follows exactly the same as those of Theorem 1 in [31], except for a new modified union bound on the detection failure event $\mathcal{G}^{c}_{k}$ derived as follows: For any $k\in\{2,\dots,N_{T}\}$,
$\mathbb{P}(\mathcal{G}^{c}_{k})$
$=\mathbb{P}\left(\left\{\exists\,l\in[k-1],\;\tau_{l}\notin\left\{\nu_{l},\dots,\nu_{l}+\ell_{l}-1\right\}\right\}\cup\left\{\tau_{k}\leq\nu_{k}\right\}\cup\mathcal{M}^{c}_{k}\cup\mathcal{L}^{c}_{k}\right)$
$\overset{(a)}{\leq}\sum_{l=1}^{k-1}\mathbb{P}\left(\tau_{l}\notin\left\{\nu_{s},\dots,\nu_{l}+\ell_{l}-1\right\},\mathcal{E}_{l-1}\right)+\mathbb{P}\left(\tau_{k}\leq\nu_{k},\mathcal{E}_{k-1}\right)+\mathbb{P}\left(\mathcal{M}^{c}_{k},\mathcal{E}_{k-1}\right)+\mathbb{P}\left(\mathcal{L}^{c}_{k},\mathcal{E}_{k-1}\right)$
$=\sum_{l=1}^{k-1}\mathbb{P}\left(\mathcal{E}_{l-1}\right)\mathbb{P}\left(\tau_{l}\notin\left\{\nu_{l},\dots,\nu_{l}+\ell_{l}-1\right\}\big|\mathcal{E}_{l-1}\right)$
$\quad\;+\mathbb{P}\left(\mathcal{E}_{k-1}\right)\mathbb{P}\left(\tau_{k}\leq\nu_{k}\big|\mathcal{E}_{k-1}\right)+\mathbb{P}\left(\mathcal{E}_{k-1}\right)\mathbb{P}\left(\mathcal{M}_{k}^{c}\big|\mathcal{E}_{k-1}\right)+\mathbb{P}\left(\mathcal{E}_{k-1}\right)\mathbb{P}\left(\mathcal{L}_{k}^{c}\big|\mathcal{E}_{k-1}\right)$
$\overset{(b)}{\leq}\sum_{l=1}^{k-1}\mathbb{P}\left(\tau_{l}\notin\left\{\nu_{l},\dots,\nu_{l}+\ell_{l}-1\right\}\big|\mathcal{E}_{l-1}\right)+\mathbb{P}\left(\tau_{k}\leq\nu_{k}\big|\mathcal{E}_{k-1}\right)+\mathbb{P}\left(\mathcal{M}_{k}^{c}\big|\mathcal{E}_{k-1}\right)+\mathbb{P}\left(\mathcal{L}_{k}^{c}\big|\mathcal{E}_{k-1}\right)$
$=\sum_{l=1}^{k}\underbrace{\mathbb{P}\left(\tau_{l}<\nu_{l}\big|\mathcal{E}_{l-1}\right)}_{\Phi_{1}}+\sum_{l=1}^{k-1}\underbrace{\mathbb{P}\left(\tau_{l}\geq\nu_{l}+\ell_{l}\big|\mathcal{E}_{l-1}\right)}_{\Phi_{2}}+\underbrace{\mathbb{P}\left(\mathcal{M}_{k}^{c}\big|\mathcal{E}_{k-1}\right)}_{\Phi_{3}}+\underbrace{\mathbb{P}\left(\mathcal{L}_{k}^{c}\big|\mathcal{E}_{k-1}\right)}_{\Phi_{4}}$ (6)
where step $(a)$ is due to the union bound, and step $(b)$ is owing to the fact that $\mathbb{P}\left(\mathcal{E}_{k-1}\right)\leq 1$. The upper bounds on $\Phi_{1}$ and $\Phi_{2}$ follows the same as those in [31]. For terms $\Phi_{3}$ and $\Phi_{4}$, we can upper bound them as follows: For any $k\in\{2,\dots,N_{T}\}$,
$\mathbb{P}\left(\mathcal{M}_{k}^{c}\big|\mathcal{E}_{k-1}\right)=$
$\mathbb{P}\Bigg(\!\exists\,l\!<\!k,c\in\mathcal{C},i\in[N_{\mathrm{e}}]:\!\sum_{t=\tau_{l-1}}^{\nu_{l}-1}\!\mathds{1}\left\{\!C_{t}=c,(t-\tau_{l-1}-1)\!\!\!\!\mod\!\!\left\lceil\frac{N_{\mathrm{e}}}{\alpha_{l}}\right\rceil\!=i-1\right\}<m_{\mathcal{D}}\Bigg|\mathcal{E}_{k-1}\!\Bigg)$
$=\mathbb{P}\Bigg(\exists c\in\mathcal{C},i\in[N_{\mathrm{e}}]:\sum_{t=\tau_{k-2}}^{\nu_{k-1}-1}\mathds{1}\{C_{t}=c,(t-\tau_{l-1}-1)\!\!\!\mod\lceil N_{\mathrm{e}}/\alpha_{l}\rceil=i-1\}<m_{\mathcal{D}}\Bigg|\mathcal{E}_{k-1}\Bigg)$
$\overset{(a)}{\leq}\sum_{c\in\mathcal{C}}\sum_{i\in[N_{e}]}\mathbb{P}\Bigg(\sum_{t=\tau_{k-2}}^{\nu_{k-1}-1}\mathds{1}\{C_{t}=c,(t-\tau_{l-1}-1)\!\!\!\mod\lceil N_{\mathrm{e}}/\alpha_{l}\rceil=i-1\}<m_{\mathcal{D}}\Bigg|\mathcal{E}_{k-1}\Bigg)$
$\overset{(b)}{\leq}\sum_{c\in\mathcal{C}}\sum_{i\in[N_{e}]}\exp\left(-2m_{\mathcal{D}}(\log T)^{2}\right)$
$\overset{(c)}{\leq}|\mathcal{C}|N_{\mathrm{e}}T^{-2}$ (7)
where step $(a)$ stems from the union bound. In step $(b)$, since $\mathds{1}\left\{C_{t}=c\right\}$ are i.i.d. $1/4$-sub-Gaussian for $t>\tau_{k-2}$ given $\mathcal{E}_{k-1}$, we can apply Chernoff bound to obtain the inequality. In step $(c)$, we simply use the fact that $m_{\mathcal{D}}>0$ and $\log T>1$ for $T\geq 3$. Note that $\mathcal{M}_{1}^{c}=\emptyset$ and thus $\Pr(\mathcal{M}_{1}^{c}|\mathcal{E}_{0})=0$. The derivation of the upper bound of $\Phi_{4}$ follows the exact same steps as in the ones in (7). The rest of the proof follows exactly the same as those in [31].
@@
In PS-PBs, $N_{\mathrm{e}}=d$, $p\geq 1/2$, and $q=r=0$. Thus, $R_{T}=\tilde{\mathcal{O}}(\sqrt{dN_{T}T}+d^{p}\sqrt{N_{T}T})=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}(\log|\Pi|)^{r}\sqrt{N_{T}T})$.
In PS-KBs, $q\geq 1/2,\ p\geq 0$ and $r=0$. We can upper bound $N_{\mathrm{e}}$ using the fact that $|\mathcal{V}_{T}|\leq\lceil\sqrt{d}R/2\delta_{T}\rceil^{d}$. Thus, $N_{e}\leq\lceil C\gamma_{T}^{2q/d}\rceil^{d}$ with $\delta_{T}=\frac{Rd^{1/2-2p/d}}{2(C\gamma_{T}^{2q})^{1/d}}$ and $R_{T}=\tilde{\mathcal{O}}((d^{2p}\gamma_{T}^{2q}N_{T}T)^{1/2}+d^{p}\gamma_{T}^{q}\sqrt{N_{T}T})=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}(\log|\Pi|)^{r}\sqrt{N_{T}T})$.
We emphasize that when the number of action is smaller than the covering number, i.e., $|\mathcal{A}|<\lceil C\gamma_{T}^{2q/d}\rceil^{d}\leq\gamma_{T}$, then we can set $\mathcal{A}_{\mathrm{e}}$ to be the entire action set $\mathcal{A}$. In this case, $N_{\mathrm{e}}<\gamma_{T}$, guaranteeing order-optimal regret.
In PS-CBs, $N_{\mathrm{e}}\leq|\mathcal{A}|$, $r\geq 1/2$, $p=q=0$, and $|\Pi|=|\mathcal{A}|^{|\mathcal{C}|}$. Thus, $R_{T}=\tilde{\mathcal{O}}((|\mathcal{A}|\log|\Pi|)^{r}\sqrt{N_{T}T}+\sqrt{|\mathcal{C}||\mathcal{A}|N_{T}T})=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}(|\mathcal{A}|\log|\Pi|)^{r}\sqrt{N_{T}T})$.